Capital rationing NPV approach r with ฮ 13 7% cost of capital must sclect the opti al group of pro ccts from those shown in the follo na table gven its capital budget of $1 10 milion. NPV at 13.7% cost of capital 585,000 14,000 27,000 Projcct Initial investment 5300,000 400,000 100,000 900,000 400,000 100,000 900,000 1,000 2,000 163,000 a. Calulate the preent vaue o cash inncws aszociated with cach project The preeent value of cseh inows for project A is sRound to the neerest doler Round to the nesresl dolr.)
Answer #1

NOTE: NPV = PV - Initial Investment OR Present Value = Initial Investment + NPV

Project Initial Investment NPV PV
A 300000 85000 (300000+85000) = $ 385000
B 400000 14000 (400000+14000) = $ 414000
C 100000 27000 (100000+27000) = $ 127000
D 900000 92000 (900000+92000) = $ 992000
E 400000 61000 (400000+61000) = $ 461000
F 100000 42000 (100000+42000) = $ 142000
G 900000 163000 (900000+163000) = $ 1063000